The volume of a cylindrical can is about 120 cubic centimeters. If the height of the can is 10 cm, determine the radius of the can. Round your answer to the nearest tenth.



Answer by mananth(16946)

The volume of a cylindrical can is about 120 cubic centimeters. If the height of the can is 10 cm, determine the radius of the can. Round your answer to the nearest tenth.
V(cylinder) = pi*r^2*h
r=radius
h=height
120 = pi*r^2*10
120/10 = pi*r^2
12=pi*r^2
12/pi= r^2
sqrt(12/pi) = r
1.9544
the radius of the cylindrical can is approximately 1.95 cm.
You round it off